dx/x(y2-z2)=dy/-y(z2+x2)=dz/z(x2+y2)


1
Expert's answer
2021-02-23T08:17:12-0500

"\frac{dx}{x(y^2-z^2)}=\frac{dy}{-y(z^2+x^2)}=\frac{dz}{z(x^2+y^2)}"

Solution:

Let's make an integral combination using the property of equal fraction:

"\frac{xdx+ydy}{x^2(y^2-z^2)-y^2(z^2+x^2)}=\frac{dz}{z(x^2+y^2)}"

"\frac{\frac12d(x^2+y^2)}{-z^2(x^2+y^2)}=\frac{dz}{z(x^2+y^2)}"

"\frac12d(x^2+y^2)=-zdz"

"\frac12(x^2+y^2)=-\frac{z^2}{2}+C"

"x^2+y^2+z^2=C_1"

"\phi_1(x,y,z)=x^2+y^2+z^2=C_1" - first integral of the system.

Another integral combination:

"\frac{ydx-xdy}{xy(y^2-z^2)+yx(z^2+x^2)}=\frac{dz}{z(x^2+y^2)}"

"\frac{y^2d\frac{x}{y}}{xy(x^2+y^2)}=\frac{dz}{z(x^2+y^2)}"

"\frac{d\frac{x}{y}}{\frac{x}{y}}=\frac{dz}{z}"

"\ln|\frac{x}{y}|=\ln|z|+\ln{C_2}"

"\frac{x}{zy}=C_2"

"\phi_2(x,y,z)=\frac{x}{zy}=C_2" - another integral of the system.

Since

rank"\begin{pmatrix}
 \frac{\partial\phi_1}{\partial x} & \frac{\partial\phi_1}{\partial y} & \frac{\partial\phi_1}{\partial z} \\
 \frac{\partial\phi_2}{\partial x} & \frac{\partial\phi_2}{\partial y} & \frac{\partial\phi_2}{\partial z}
\end{pmatrix}=
\begin{pmatrix}
 2x & 2y & 2z \\
 \frac{1}{zy} & - \frac{x}{zy^2} & -\frac{x}{yz^2}
\end{pmatrix}=2" ,

then "\phi_1" and "\phi_2" are independent.

General solution:

"u=\varPhi(\phi_1, \phi_2)=\varPhi(x^2+y^2+z^2, \frac{x}{yz})" , where "\varPhi" is any continues differentiable function of arguments "\phi_1" and "\phi_2" .

Answer: "u=\varPhi(x^2+y^2+z^2, \frac{x}{yz})" .